OpenStudy (anonymous):

Find the limit of \[\mathop {\lim }\limits_{x \to \infty } {\left( {\frac{{\sqrt[x]{a} + \sqrt[x]{b}}} {2}} \right)^x}\]

OpenStudy (anonymous):

Today is suck...!!

OpenStudy (anonymous):

\[\Huge{\lim_{x \rightarrow \infty}e^{x \:\:log_e (\frac{ a^{\frac{1}{x}} + b^{\frac{1}{x}}}{2})}}\]

OpenStudy (anonymous):

ah Lol sorry I am getting no where

OpenStudy (anonymous):

ah wait I got it

OpenStudy (anonymous):

\[\Huge{\lim_{x \rightarrow \infty}e^{x \:\:log_e (\frac{ a^{\frac{1}{x}} + b^{\frac{1}{x}}}{2})}}\] Now \(\frac{1}{x} \rightarrow 0\) \[\Huge{\lim_{x \rightarrow \infty}e^{x \:\:log_e (\frac{ a^0 + b^0}{2})}}\]

OpenStudy (anonymous):

\[\Huge{\lim_{x \rightarrow \infty}e^{x \:\:log_e 1}}\]\[log_e = 0 \] \[e^0= 1 \] ah It must be the answer I can be wrong too

OpenStudy (anonymous):

i think u r correct....!!!!!!

OpenStudy (anonymous):

this step is alright. I have done this \[\mathop {\lim }\limits_{x \to \infty } {e^{x\ln \frac{{\sqrt[x]{a} + \sqrt[x]{b}}} {2}}}\] is right But \[\mathop {\lim }\limits_{x \to \infty } x\ln \frac{{\sqrt[x]{a} + \sqrt[x]{b}}} {2}\] is not zero because x is infinit, so here it's time to use l'hospital rule. But if you do next you will see it become messy. By the way, how to increase the size of the letter of equation. ? mine is too small to see.

OpenStudy (anonymous):

I can be wrong

OpenStudy (anonymous):

\Huge{--------}

OpenStudy (anonymous):

You can increase the size through that

OpenStudy (anonymous):

Usually, I use l'hospital rule, But this time it becomes complex and don't work.

OpenStudy (anonymous):

I mean you get stuck sometimes very bad ....using l'hospital's rule

OpenStudy (anonymous):

I like to solve Limits through natural way you should try that too it's so cool to solve them without l'Hospital's rule

OpenStudy (anonymous):

I agree. It is the time it doesn't work. ╮(╯▽╰)╭

OpenStudy (anonymous):

the answer from textbook is \[\sqrt {ab} \] I am trying to solve it.

OpenStudy (anonymous):

ah I might be wrong I will solve this one ...again and see what wrong I did

OpenStudy (anonymous):

Good luck for both of us..

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=lim+%28%28a^%281%2Fx%29%2Bb^%281%2Fx%29%29%2F%282%29%29^x++++++++as+x-%3Einfinity+ I tried but I am unable to get the answer and I was wrong earlier

OpenStudy (anonymous):

I use l'hospital rule and finally get the answer. I need substitution.

OpenStudy (anonymous):

wow... the link you give me is very powerful... so cool... Thanks so much.